OpenStudy (anonymous):

What formula would i use to get the answer to this question ? A musician is planning to market a CD. The fixed costs are $980 and the variable costs are $4 per CD. The wholesale price of the CD will be $11. For the artist to make a profit, revenues must be greater than costs. How many CDs, x, must be sold for the musician to make a profit?

OpenStudy (anonymous):

What is the relation between benefit and revenue?

OpenStudy (anonymous):

@aanders Can you reply?

OpenStudy (anonymous):

definition wise?

OpenStudy (anonymous):

\[ (11 x-(980 + 4 x ))>0 \implies x>140\] Hence a minimum of value of x to make a profit is 141.

OpenStudy (anonymous):

or in other words 141 CDS must be sold for the musician to make a profit

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Well done!

OpenStudy (anonymous):

The cost for the production of x CDs is given by the equation c(x)=980+4*x The money gathered by selling x CDs in given by w(x) = 11*x we know that profit is the wholesale price minus the production cost, or mathematically: p(x) = w(x) - c(x) so, when profit is zero, we get the minimum amount of CDs to avoid loss. ...foolformath beat me to it

OpenStudy (anonymous):

thank you both though for the explanations!